Suppose that the economy of Stormwind has the following​ features:

The level of private saving is ​$200.

The level of investment expenditure is ​$100.

The level of exports is ​$75.

The level of imports is ​$125.

If​ Stormwind's economy is in​ equilibrium, the level of public saving would be equal to ​$

nothing


. ​(Round your response to the nearest​ dollar.)

The level of public saving shows that the government in Stormwind has a budget ▼

 

deficit

surplus

.

If the level of government purchases is ​$200 in​ Stormwind, the level of taxes must be ​$

nothing


. ​(Round your response to the nearest​ dollar.)

1
Expert's answer
2021-02-16T10:06:50-0500

Private saving=spr=Y-T-C

public saving =sp=T-G

In equilibrium I=S

S=SPR+SP

200=100+SP

SP=$100

the level of public saving shows that the government If stormwind has a surplus budget.

If the government purchases $200 the level of taxes must be $100